# London WC2R 2LS, UK

### v1

@@ 4.3.2 HKT geometry and conformal invariance
Next suppose that $M^{4k}$ is an HKT manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$ – again $\widehat{V}$ is normalised to have length $1$. As for KT manifolds, the conditions ^$\widehat{\nabla}\widehat{V}=0$ and ^$\widehat{\nabla}I_{r}=0$ imply that ^$\widehat{\nabla}\widehat{Z}_{r}=0$, where $\widehat{Z}_{r}^{J}=-I_{r}{}^{J}{}_{K}\widehat{V}^{K}$. The vector fields $\widehat{V}$ and $\widehat{Z}_{r}$ are orthogonal $G(\widehat{V},\widehat{Z}_{r})=0$, $G(\widehat{Z}_{r},\widehat{Z}_{s})=0$ for $r\not=s$, and $\widehat{Z}_{r}$ have length $1$ as well. Furthermore, it can be shown that $\widehat{Z}_{r}$, as $\widehat{V}$, leave invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Note that $\iota_{\widehat{Z}_{r}}F=0$ as a consequence of $\iota_{\widehat{V}}F=0$ and that $F$ is an (1,1)-form with respect to each $I_{r}$. The proof that $H$ and $F$ are invariant under the action of $\widehat{V}$ and $\widehat{Z}_{r}$ is similar to that given in (101) and (98), respectively. One consequence of ^$\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}_{r}=0$ is that the holonomy of $\widehat{\nabla}$ reduces to a subgroup of $Sp(k-1)$.
To continue, let $\mathfrak{g}$ be the Lie algebra spanned by $\widehat{V}$ and $\widehat{Z}_{r}$ as well as all their commutators including the vector fields constructed by acting with $I_{r}$ on the commutators. All these vector fields, denoted by $\widehat{V}_{a}$, will be $\widehat{\nabla}$ covariantly constant, $\iota_{\widehat{V}_{a}}F=0$ and leave invariant $G$, $H$ and $F$. To proof of the invariance of $G$, $H$ and $F$ follows from the arguments presented in the KT case and in particular those in (101), (98), (102), (103 and (105). As in the KT case the metric and torsion can be written as in (110) but now $\mathfrak{g}$ has dimension $4q$ and the holonomy of $\widehat{\nabla}$ reduces to $Sp(k-q)$. Furthermore, the Hermitian forms can be written as
@@
To construct examples of HKT geometries that solve the conditions for conformal invariance consider principal bundles with fibre a group manifold $\mathcal{G}$ with a left invariant HKT structure equipped with a connection $\lambda$ over a base space $N^{4k-4q}$ whose geometry will be described later. The left invariant HKT structure on $\mathcal{G}$ with metric $G^{\mathcal{G}}$, torsion $H^{\mathcal{G}}$ and Hermitian forms $I_{r}^{\mathcal{G}}$ is
@@
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of ${\mathcal{G}}$ evaluated along the vertical directions, i.e. $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}\lambda^{\alpha}$ instead of $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}L^{\alpha}$, and $\widehat{\theta}^{\perp}$ is the Lee form of $N^{4k-4q}$. As $\widehat{\theta}^{\mathcal{G}}$ is $\widehat{\nabla}$-covariantly constant, the condition for conformal invariance becomes
^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~{},$ (144)
i.e. it reduces to demanding that the base space $N^{4k-4q}$ satisfies the condition for conformal invariance. This can be solved by taking $N^{4k-4q}$ to be conformally balanced.
@@
Thus one can write $(B_{\alpha})_{rs}=B^{t}_{\alpha}\epsilon_{trs}$. Furthermore, taking the Lie derivative of (145) with respect to $\mathcal{L}_{\widehat{V}_{\beta}}$ and using $[\mathcal{L}_{\widehat{V}_{\alpha}},\mathcal{L}_{\widehat{V}_{\beta}}]=\mathcal{L}_{[\widehat{V}_{\alpha},\widehat{V}_{\beta}]}$, we find that
$[B_{\alpha},B_{\beta}]^{r}{}_{s}=-H_{\alpha\beta}{}^{\gamma}(B_{\gamma})^{r}{}_{s}~{},$ (147)
@@
where the index $r$ should not be summed over. These are additional conditions on the representation $(B_{\alpha})$ of $\mathfrak{g}$. The last condition in (124) is independent from the remaining two and has to be satisfied as well. It imposes the integrability of $I_{r}$ on the base space $N^{4k-4q}$. It remains to investigate the properties of the connection $\lambda$ and to determine the geometry of the base space. For the former, notice from (145) that
@@
^$\widehat{\nabla}_{\mu}I_{rij}$ $=\partial_{\mu}I_{rij}-\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{i}I_{rkj}+\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{j}I_{rki}-\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{i}I_{rkj}+\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{j}I_{rki}$ (151)
=^$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}+\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{i}I_{rkj}-\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{j}I_{rki}$ (152)
=^$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}-\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}I_{sij}=0~{}.$ (153)
where ^$\widehat{\nabla}_{\mu}^{\perp}$ is the frame connection of $N^{4k-4q}$ with torsion $H^{\perp}$. Therefore, $N^{4k-4q}$ admits a Quaternionic-Kähler with torsion (QKT) structure [^18] A QKT manifold admits locally three complex structures $I_{r}$ that satisfy the algebra of imaginary unit quaternions and the covariant constancy condition ^$\widehat{\nabla}I_{r}-Q^{s}{}_{r}I_{t}=0$, where $\widehat{\nabla}$ is the metric connection with torsion a 3-form $H$ and $Q$ is an $\mathfrak{sp}(1)$ connection. which has been investigated before in HPO . Reversing the construction one can start from a QKT manifold $N^{4k-4q}$ with torsion a $(2,1)\oplus(1,2)$ form with respect to all complex structures $I_{r}^{\perp}$ such that
^$\widehat{\nabla}_{\mu}^{\perp}I_{rij}-Q_{\mu}{}^{s}{}_{r}I_{sij}=0~{}.$ (154)
Then choose an HKT group manifold $\mathcal{G}$ that admits a 3-dimensional orthogonal representation $(B_{\alpha})$ that satisfies the conditions (148). Next construct a principal bundle $M^{4k}$ with fibre $\mathcal{G}$ that admits a connection $\lambda$ that is a $\mathfrak{sp}(k)\oplus\mathfrak{sp}(1)$ instanton and set $\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}=Q_{\mu}{}^{s}{}_{r}$. This will solve all the conditions for $M^{4k}$ to admit a HKT structure with $\widehat{\nabla}$-covariantly constant vector fields generated by the $\mathcal{G}$ group action. The only remaining condition is to impose (138) on $H^{\perp}$. To investigate whether such manifolds satisfy the condition for conformal invariant, one has to compute the Lee form. It turns out that such manifolds satisfy the conformal condition provided that (144) holds. This can be satisfied for conformally balanced QKT manifold $N^{4k-4q}$.
@@ 4.3.3 $G_{2}$ and $\mathrm{Spin}(7)$ geometry and conformal invariance
If the connection $\widehat{\nabla}$ on $M^{7}$ has $G_{2}$ holonomy, the existence of an additional $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $SU(3)$. This is because the isotropy group of a vector in the 7-dimensional representation of $G_{2}$ is $SU(3)$. The $SU(3)$ structure is associated with the (almost) “Hermitian” form $\iota_{\widehat{V}}\varphi$ and the $(3,0)\oplus(0,3)$ form $\iota_{\widehat{V}}{}^{*}\varphi$, where $\varphi$ is the $\widehat{\nabla}$-covariantly constant form $G_{2}$ on $M^{7}$. Of course both $\iota_{\widehat{V}}\varphi$ and $\iota_{\widehat{V}}{}^{*}\varphi$ are $\widehat{\nabla}$-covariantly constant. Examples can be constructed as circle principal fibrations over CYT 6-dimensional manifolds $N^{6}$ with a principal bundle connection $\lambda$ an $SU(3)$ instanton. The metric and torsion on $M^{7}$ can be constructed as in (110), where $\alpha$ takes value $\alpha=0$. Moreover, the $G_{2}$ form on $M^{7}$ is
@@
where $I^{\perp}$ and $\chi^{\perp}$ are the Hermitian and $(3,0)\oplus(0,3)$ of the CYT manifold $M^{6}$ pulled back on $M^{7}$. One can demonstrate that $\mathcal{L}_{\widehat{V}^{0}}\varphi=0$. The Lee form $\widehat{\theta}=\widehat{\theta}^{\perp}$, where $\widehat{\theta}^{\perp}$ is the Lee form of $N^{6}$. Thus the condition for conformal invariance becomes ^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2d\Phi)=0$, Therefore the condition for conformal invariance can also be satisfied for $N^{6}$ a conformally balanced CYT manifold.
Similarly, if the connection $\widehat{\nabla}$ on $M^{8}$ has $\mathrm{Spin}(7)$ holonomy, the existence of an additional $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $G_{2}$. This is because the isotropy group of a vector in the 8-dimensional spinor representation of $\mathrm{Spin}(7)$ is $G_{2}$. The $G_{2}$ structure on $M^{8}$ is associated with the 3-form $\iota_{\widehat{V}}\phi$, where $\phi$ is the $\widehat{\nabla}$-covariantly constant self-dual form on $M^{8}$ associated with the $\mathrm{Spin}(7)$ structure. Clearly, $\iota_{\widehat{V}}\phi$ is also $\widehat{\nabla}$-covariantly constant. as in the $G_{2}$ case above, examples can be constructed as circle principal fibrations over conformally balanced $G_{2}$ 7-dimensional manifolds $N^{7}$ with torsion equipped with a connection $\lambda^{0}$, $(\alpha=0)$, a $G_{2}$ instanton. The metric $G$ and torsion $H$ on $M^{8}$ can be constructed as in (110). The self-dual 4-form on $M^{8}$ can be written as

### v2

@@ 4.3.3 HKT geometry and conformal invariance
Next suppose that $M^{4k}$ is an HKT manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$ – again $\widehat{V}$ is normalised to have length $1$. As for KT manifolds, the conditions $\widehat{\nabla}\widehat{V}=0$ and $\widehat{\nabla}I_{r}=0$ imply that $\widehat{\nabla}\widehat{Z}_{r}=0$, where $\widehat{Z}_{r}^{J}=-I_{r}{}^{J}{}_{K}\widehat{V}^{K}$. The vector fields $\widehat{V}$ and $\widehat{Z}_{r}$ are orthogonal $G(\widehat{V},\widehat{Z}_{r})=0$, $G(\widehat{Z}_{r},\widehat{Z}_{s})=0$ for $r\not=s$, and $\widehat{Z}_{r}$ have length $1$ as well. Furthermore, it can be shown that $\widehat{Z}_{r}$, as $\widehat{V}$, leave invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Note that $\iota_{\widehat{Z}_{r}}F=0$ as a consequence of $\iota_{\widehat{V}}F=0$ and that $F$ is an (1,1)-form with respect to each $I_{r}$. The proof that $H$ and $F$ are invariant under the action of $\widehat{V}$ and $\widehat{Z}_{r}$ is similar to that given in (101) and (98), respectively. One consequence of $\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}_{r}=0$ is that the holonomy of $\widehat{\nabla}$ reduces to a subgroup of Sp(k1)Sp(k-1).
To continue, let $\mathfrak{g}$ be the Lie algebra spanned by $\widehat{V}$ and $\widehat{Z}_{r}$ as well as all their commutators including the vector fields constructed by acting with $I_{r}$ on the commutators. All these vector fields, denoted by $\widehat{V}_{a}$, will be $\widehat{\nabla}$ covariantly constant, $\iota_{\widehat{V}_{a}}F=0$ and leave invariant $G$, $H$ and $F$. To proof of the invariance of $G$, $H$ and $F$ follows from the arguments presented in the KT case and in particular those in (101), (98), (102), (103 and (105). As in the KT case the metric and torsion can be written as in (110) but now $\mathfrak{g}$ has dimension $4q$ and the holonomy of $\widehat{\nabla}$ reduces to Sp(kq)Sp(k-q). Furthermore, the Hermitian forms can be written as
@@
To construct HKT geometries that solve some of the conditions for conformal invariance consider principal bundles with fibre a group manifold $\mathcal{G}$ with a left invariant HKT structure equipped with a connection $\lambda$ over a base space $N^{4k-4q}$ whose geometry will be described later. The left invariant HKT structure on $\mathcal{G}$ with metric $G^{\mathcal{G}}$, torsion $H^{\mathcal{G}}$ and Hermitian forms $I_{r}^{\mathcal{G}}$ is
@@
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of $\mathcal{G}$} evaluated along the vertical directions, i.e. $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}\lambda^{\alpha}$ instead of $\widehat{\theta}^{\mathcal{G}}=\widehat{\theta}^{\mathcal{G}}_{\alpha}L^{\alpha}$, and $\widehat{\theta}^{\perp}$ is the Lee form of $N^{4k-4q}$. As $\widehat{\theta}^{\mathcal{G}}$ is $^\widehat{\nabla}$-covariantly constant, the condition for conformal invariance becomes
$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~,$ (144)
i.e. it reduces to demanding that the base space $N^{4k-4q}$ satisfies the condition for conformal invariance. This can be solved by taking $N^{4k-4q}$ for example to be conformally balanced, $\widehat{\theta}^{\perp}=-2{\mathrm{d}}\Phi$ or hyper-Kähler with $\widehat{\theta}^{\perp}=0$ and $\Phi$ constant.
@@
Thus one can write $(B_{\alpha})_{rs}=B^{t}_{\alpha}\epsilon_{trs}$. Furthermore, taking the Lie derivative of (145) with respect to $\mathcal{L}_{\widehat{V}_{\beta}}$ and using $\mathcal{L}_{\widehat{V}_{\alpha}},\mathcal{L}_{\widehat{V}_{\beta}}]=\mathcal{L}_{[\widehat{V}_{\alpha},\widehat{V}_{\beta}]}$, we find that
$B_{\alpha},B_{\beta}]^{r}{}_{s}=-H_{\alpha\beta}{}^{\gamma}(B_{\gamma})^{r}{}_{s}~$, (147)
@@
where the index $r$ should not be summed over. These are additional conditions on the representation $(B_{\alpha})$ of $\mathfrak{g}$. The last condition in (124) is independent from the remaining two and has to be satisfied as well. It imposes the integrability of $I_{r}$ on the base space $N^{4k-4q}$.
It remains to investigate the properties of the connection $\lambda$ and to determine the geometry of the base space. For the former, notice from (145) that
@@
$\widehat{\nabla}_{\mu}I_{rij}$ $=\partial_{\mu}I_{rij}-\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{i}I_{rkj}+\mathbf{e}_{\mu}^{m}\widehat{\Omega}_{m}{}^{k}{}_{j}I_{rki}-\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{i}I_{rkj}+\lambda_{\mu}^{\alpha}\widehat{\Omega}_{\alpha}{}^{k}{}_{j}I_{rki}$ (151)
$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}+\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{i}I_{rkj}-\lambda_{\mu}^{\alpha}\mathcal{F}_{\alpha}{}^{k}{}_{j}I_{rki}$ (152)
$=\widehat{\nabla}_{\mu}^{\perp}I_{rij}-\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}I_{sij}=0~.$ (153)
where $\widehat{\nabla}_{\mu}^{\perp}$ is the frame connection of $N^{4k-4q}$ with torsion $H^{\perp}$. Therefore, $N^{4k-4q}$ admits a Quaternionic-Kähler with torsion (QKT) structure [^21] A QKT manifold admits locally three complex structures $I_{r}$ that satisfy the algebra of imaginary unit quaternions and the covariant constancy condition $\widehat{\nabla}I_{r}-Q^{s}{}_{r}I_{t}=0$, where $\widehat{\nabla}$ is the metric connection with torsion a 3-form $H$ and $Q$ is an $\mathfrak{sp}(1)$ connection. which has been investigated before in HPO . Reversing the construction one can start from a QKT manifold $N^{4k-4q}$ with torsion a $(2,1)\oplus(1,2)$ form with respect to all complex structures $I_{r}^{\perp}$ such that
$\widehat{\nabla}_{\mu}^{\perp}I_{rij}-Q_{\mu}{}^{s}{}_{r}I_{sij}=0~.$ (154)
Then choose an HKT group manifold $\mathcal{G}$ that admits a 3-dimensional orthogonal representation $(B_{\alpha})$ that satisfies the conditions (148). Next construct a principal bundle $M^{4k}$ with fibre $\mathcal{G}$ that admits a connection $\lambda$ that is a $\mathfrak{sp}(k)\oplus\mathfrak{sp}(1)$ instanton and set $\lambda_{\mu}^{\alpha}(B_{\alpha})^{s}{}_{r}=Q_{\mu}{}^{s}{}_{r}$. This will solve all the conditions for $M^{4k}$ to admit a HKT structure with $^\widehat{\nabla}$-covariantly constant vector fields generated by the $\mathcal{G}$ group action.
To investigate whether such manifolds satisfy the conditions for conformal invariance in (93), one has to compute the Lee form. It turns out that such manifolds satisfy this conformal condition provided that (144) holds, i.e. it is required that $(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)$ is $\widehat{\nabla}^{\perp}$-covariantly constant. This for example can be satisfied by a conformally balanced QKT manifold $N^{4k-4q}$. The only remaining condition is to impose (138) on $H^{\perp}$. This is rather involved in general, especially for examples for which 4k4q>44k-4q>4, and results will be reported elsewhere.
@@ 4.3.4 $G_{2}$ and $\mathrm{Spin}(7)$ geometry and conformal invariance
If the connection $\widehat{\nabla}$ on $M^{7}$ has $G_{2}$ holonomy, the existence of an additional $^\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $SU(3)$. This is because the isotropy group of a vector in the 7-dimensional representation of $G_{2}$ is $SU(3)$. The $SU(3)$ structure is associated with the (almost) “Hermitian” form $\iota_{\widehat{V}}\varphi$ and the $(3,0)\oplus(0,3)$ form $\iota_{\widehat{V}}{}^{*}\varphi$, where $\varphi$ is the $^\widehat{\nabla}$-covariantly constant form $G_{2}$ on $M^{7}$. Of course both $\iota_{\widehat{V}}\varphi$ and $\iota_{\widehat{V}}{}^{*}\varphi$ are $^\widehat{\nabla}$-covariantly constant.
Therefore to solve some of the geometric conditions for conformal invariance, one can begin from circle principal fibrations over CYT 6-dimensional manifolds $N^{6}$ with a principal bundle connection $\lambda$ an $SU(3)$ instanton. The metric and torsion on $M^{7}$ can be constructed as in (110), where $\alpha$ takes value $\alpha=0$. Moreover, the $G_{2}$ form on $M^{7}$ is
@@
where $I^{\perp}$ and $\chi^{\perp}$ are the Hermitian and $(3,0)\oplus(0,3)$ of the CYT manifold $M^{6}$ pulled back on $M^{7}$. One can demonstrate that $\mathcal{L}_{\widehat{V}^{0}}\varphi=0$. The Lee form $\widehat{\theta}=\widehat{\theta}^{\perp}$, where $\widehat{\theta}^{\perp}$ is the Lee form of $N^{6}$. Thus the condition for conformal invariance becomes $\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2d\Phi)=0$, So, this condition for conformal invariance can be satisfied for $N^{6}$ a conformally balanced CYT manifold $\widehat{\theta}^{\perp}=-2d\Phi$.
Similarly, if the connection $\widehat{\nabla}$ on $M^{8}$ has $\mathrm{Spin}(7)$ holonomy, the existence of an additional $^\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ will reduce the holonomy to a subgroup of $G_{2}$. This is because the isotropy group of a vector in the 8-dimensional spinor representation of $\mathrm{Spin}(7)$ is $G_{2}$. The $G_{2}$ structure on $M^{8}$ is associated with the 3-form $\iota_{\widehat{V}}\phi$, where $\phi$ is the $^\widehat{\nabla}$-covariantly constant self-dual form on $M^{8}$ associated with the $\mathrm{Spin}(7)$ structure. Clearly, $\iota_{\widehat{V}}\phi$ is also $^\widehat{\nabla}$-covariantly constant.
As in the $G_{2}$ case above to solve some of the conditions for conformal invariance, one can begin from circle principal fibrations over conformally balanced $G_{2}$ 7-dimensional manifolds $N^{7}$ with torsion equipped with a connection $\lambda^{0}$, $(\alpha=0)$, a $G_{2}$ instanton. The metric $G$ and torsion $H$ on $M^{8}$ can be constructed as in (110). The self-dual 4-form on $M^{8}$ can be written as